OpenStudy (anonymous):

How long will it take a population to triple if it is growing at a rate of 2.4% continuously?

OpenStudy (zzr0ck3r):

what is the equation for continuous growth?

OpenStudy (zzr0ck3r):

hint: pert

OpenStudy (dan815):

xD

OpenStudy (anonymous):

e^0.024t=3 ? is that the equation?

OpenStudy (zzr0ck3r):

well we would have \(A=Pe^{rt} \) but now we want \(Pe^{rt}=3P\)

OpenStudy (zzr0ck3r):

so \(e^{0.024*t}=3\)

OpenStudy (zzr0ck3r):

can you solve for t?

OpenStudy (anonymous):

You would take the log of e^(0.024)(t) and log of 3?

OpenStudy (anonymous):

then divide the log of 4 by o.o24?

OpenStudy (anonymous):

log of 3*

OpenStudy (zzr0ck3r):

try this \(ln\) of both sids \(ln(e^{0.024t})=ln(3)\\ 0.024t*ln(e)=ln(3)\\0.24t*1=ln(3)\\t=\frac{ln(3)}{0.024}\)

OpenStudy (anonymous):

20 years?

OpenStudy (zzr0ck3r):

if you take ln of one side you must do ln of the other, you cant do natural log on one side and log base 3 on the other

OpenStudy (zzr0ck3r):

if thats what that is....

OpenStudy (zzr0ck3r):

I aint no calculator:)

OpenStudy (anonymous):

Lol :D Let me do it real quick. Pretty sure I did Log not Ln

OpenStudy (anonymous):

Thanks so much ^_^

OpenStudy (zzr0ck3r):

np